OpenStudy (anonymous):

If : \(\color{blue}{\text{2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38 = 200}}\), Then, \(\color{red}{\text{Add}}\) any \(\color{cyan}{\text{Five}}\) numbers and get \(\color{green}{100}\)..

OpenStudy (aravindg):

Colorful question :)

OpenStudy (e.cociuba):

How do u do that? I have been trying to figure it out lol

OpenStudy (johnweldon1993):

\(\color{red}{\text{What?}}\) \(\color{blue}{\text{Make it colorful?}}\)

OpenStudy (e.cociuba):

Haha ya tht^^ :P

OpenStudy (johnweldon1993):

And @annanya I don't believe it's possible...

OpenStudy (johnweldon1993):

\(\color{green}{\text{e.cociuba}}\) \(\color{lightblue}{\text{It's easy :P}}\)

OpenStudy (jhannybean):

hint: right click on @johnweldon1993's text and click "show math as" "tex commands" :P

OpenStudy (e.cociuba):

mmhhm says the guy who can do it lol

OpenStudy (johnweldon1993):

NO!!! lol WHY @Jhannybean !!! haha

OpenStudy (anonymous):

I presume you are using the word 'sum' in the more mathematical sense of 'added total' rather than the everyday sense of 'calculation,. None of the numbers in your list is divisible by 4. In fact, mod 4, each leaves remainder 2, ie. each can be expressed as 4n+2 for some integer n. Therefore the total can be expressed as: 4a+2 + 4b+2 + 4c+2 + 4d+2 + 4e+2 = 4(a+b+c+d+e+2) + 2 Since 100 is a multiple of 4, this total cannot equal 100. I hope this helps. P.S. If you just mean 'sum' in the sense of 'calculation' then 2x6x14-38-30 = 100

OpenStudy (jhannybean):

xD